A boat is moving at a speed of 100 km/hr.
 
Suddenly, the still water beneath it starts flowing at the rate of 20 km/hr.
 
If the coefficient of friction is 0.001 and water current is flowing in opposite direction, the net speed is X.
 
If the coefficient of friction is 0.001 and water current is flowing in same direction as the boat, the net speed is Y.
 
The original speed is Z.
 
Then, what is the mass of the man who is driving the boat?
 
And isn’t Ronaldo still the greatest of all times?

Grade:11

1 Answers

Arun
3 years ago
Since the boat moves across the river, the resultant velocity of the boat is :
 
60² + 25² = 4225
 
Rv = √4225 = 65
 
= 65km/h
 
The speed = 65 km/h
 
The distance travelled = 195km.
 
Time = 195/65 = 3 hrs
 
Assuming the boat crossed straight across the river in still waters :
 
Speed = 60 km/h
 
Time = 3 hrs
 
Width of the river :
 
= 60 × 3 = 180km
 
= 180km